OpenStudy (anonymous):

Gunpowder can be prepared by saltpetre and nitrous oxide. Price of saltpetre is thrice the price of nitrous oxide. Notorious gangster pitbull sells the gunpowder at $2160 per 10 g, thereby making a profit of 20% . if the ratio of saltpetre and nitrous oxide in the mixture be 2:3, find the cost price of saltpetre.

OpenStudy (anonymous):

help!!

OpenStudy (anonymous):

I am thinking, lol

OpenStudy (anonymous):

in this question they are through too much information.. I THINK

OpenStudy (anonymous):

the ans is not given too

OpenStudy (anonymous):

OK I think i got it..

OpenStudy (anonymous):

Have u got it??

OpenStudy (anonymous):

So now, we know that notorious gangster pitbull sells the gunpowder at $2160 per 10 g... Therefore for 1 gram it is $216 we know that the price of saltpetre is thrice the price of nitrous oxide.. so if the price of nitrous oxide is x, then saltpetre is 3x and for one gram 3x+x=216 4x= 216; and x= $54 therefore for ONE gram of this gun powder it is $162 for the salt and 54 for the nitrous. Since we know that the ratio of salt to nitrous in the powder is 2:3 then 2/5 *10 grams = 4 grams of salt and 3/5*10 grams =6 grams of nitrous therefore to find the total price of saltpetre in the 10 gram solution: 4*162= $648 since pitbull makes 20% profit from that 648 dollars, the cost price of saltpetre is 648 - 648 *0.2 = 518.4 DOLLARS is the cost price.

OpenStudy (anonymous):

let me know what you think

OpenStudy (anonymous):

thats correct..

OpenStudy (anonymous):

what you got the final answer?

OpenStudy (anonymous):

or you think it is correct? lol

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok cool ..

OpenStudy (anonymous):

i think its correct